# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v1

@@ Abstract
It is worth noting that implicit behind the picture is the separation between two time scales. One is the time scale for local equilibration among wave packets, which drives the distribution towards $f_{0}(\tilde{\varepsilon}_{l})$. It is associated with the relaxation of the local energy shift $\delta\varepsilon_{l}$ (This is also the process that establishes $f_{0}(\varepsilon_{l})$ in the absence of field). The other is the time scale for the macroscopic transport, which might be regarded as a relaxation of the electric potential energy. The latter is nonlocal in the sense that the relaxation requires the transport of electron across a macroscopic distance (through the sample). Clearly, in a diffusive system, the former time scale (on the order of $\tau$) is much shorter than the latter. This justifies the use of $f_{0}(\tilde{\varepsilon}_{l})$ as a starting point for studying transport process.
Figure 1: (a) Crystal structure of orthorhombic CuMnAs. The local moment direction in the antiferromagnetic ground state is marked. (b) Calculated band structure. (c) Calculated nonlinear response $\chi^{(2\omega)}_{xxx}$ versus chemical potential $\mu$. Temperature is set at 303030~{}30K. (d) Distribution of BCP dipole (the summand in Eq. (19) in the $k_{y}=0$ plane in momentum space. The dashed line marks the nodal loop in the band structure around $X$ point, which is slightly gapped under spin-orbit coupling.
Finally, we note that the frequency dependence of $j^{(2\omega)}$ under ac driving [Eq. (17) versus (18)] is a result of this new understanding, which can be probed in experiment. Experimentally, one typically applies an ac driving at frequency $\omega$ and measures the double-frequency response using lock-in technique [10, 11]. From our results, one can see that at low frequencies $\omega\ll\tau^{-1}$, the $2\omega$ response is purely Hall (dissipationless). If $\tau$ is sufficiently large that one can tune $\omega$ into the regime $\tau^{-1}\ll\omega<\Delta$, then there will also appear a non-Hall (dissipative) component in the transverse double-frequency response. This can be distinguished in experiment by the sum-frequency generation method developed by Gao et al. [29]. Meanwhile, in the high-frequency regime, there also emerges a double-frequency component in longitudinal response. The corresponding nonlinear conductivity is
$\chi_{aaa}^{(2\omega)}=-\sum_{l}V_{a}^{ll}G_{aa}^{l}f^{\prime}_{0}(\varepsilon_{l})/2.$ (19)
To have an estimation of this quantity, we consider the example of orthorhombic CuMnAs, which is an antiferromagnet with Néel temperature $T_{\mathrm{N}}\approx 300$ K [55, 56, 57, 58, 59, 60, 61]. Previous experiments [58, 59] reported that the Néel vector is aligned along the $b$ axis [Fig. 1(a)]. The magnetic band structure obtained from first-principles calculations (details in [46]) is shown in Fig. 1(b), which agrees with previous calculations [58]. This material belongs to the $m^{\prime}mm$ magnetic point group. The symmetry-allowed nonlinear longitudinal response is $\chi_{xxx}$. The result calculated from Eq. (19) is shown in Fig. 1(c). Interestingly, we find that the sizable $\chi_{xxx}$ has a large contribution from a nodal loop feature near the Fermi level [see Fig. 1(d)]. At $T=30~{}$K, the value of $\chi_{xxx}$ can reach $\sim 2.3~{}$mA///V2. From previous experiment [59], $\tau\sim 1$ ps at 30 K, and from the band structure, $\Delta\sim 20$ meV ($30$ THz). Hence, this scattering-independent nonlinear response $\chi_{xxx}$ can be observed in the frequency window $\mathrm{0.5~{}THz}\ll\omega<\mathrm{30~{}THz}$. This implies that orthorhombic CuMnAs is a promising candidate for terahertz isolation [62, 63, 64] of intrinsic nonlinear transport.

### v2

@@ Abstract
IAD arises from the local energy shift of electron wave packets. Implicit behind this picture is the separation between two time scales. One is the time scale for local equilibration among wave packets, which drives the distribution towards $f_{0}(\tilde{\varepsilon}_{l})$. It is associated with the relaxation of the local energy shift $\delta\varepsilon_{l}$ (This is also the process that establishes $f_{0}(\varepsilon_{l})$ in the absence of field). The other is the time scale for macroscopic transport, which might be regarded as a relaxation of electric potential energy. The latter is nonlocal in the sense that the relaxation requires the transport of electron across a macroscopic distance (through the sample). Clearly, in a diffusive system, the former time scale is much faster than the latter, which justifies the use of $f_{0}(\tilde{\varepsilon}_{l})$ as a starting point for studying transport process.
We point out that the notion of IAD is universal, and its importance is not limited to nonlinear transport effects. It should be associated with any uniform energy shift for diffusive systems in steady state. Actually, we find that failing to incorporate IAD will lead to violation of fundamental physics relations or principles. For example, it is known that on the basis of Einstein relation, the current response to $E$ field can be connected to that to chemical-potential gradient. The intrinsic nonlinear current driven by chemical-potential gradient was reported in Ref. [55]. In End Matter, we show that it is Eq. (7) from (D3), rather than Eq. (6), that fulfills the connection. In other words, IAD must be incorporated to be consistent with Einstein relation. As a second example, it is well known that to linear order in a uniform magnetic field $B$, the electrons acquire a local energy shift $\delta\varepsilon_{l}=-\bm{m}_{l}\cdot\bm{B}$, where $\bm{m}_{l}$ is the spin plus orbital magnetic moment of the wave packet [56, 5]. In the same spirit, this shift also induces an IAD, which must be considered. One can easily show that failing to account for this IAD would lead to a non-vanishing transport current driven a uniform dc $B$ field (see End Matter), which violates fundamental physics principles.
Our discussion of IAD here is using the semiclassical theory framework, nevertheless, the revealed physics should be general. This is because (1) it enjoys a transparent and coherent physical picture, and (2) the above evidences demonstrate it is indispensable for any theory that wants to maintain consistency with fundamental physical relations/principles.
We also mention that although phrased using familiar ‘classical’ concepts, the semiclassical theory is a quantum theory, which fully captures quantum effects (including interband coherence) in the intraband transport regime. Meanwhile, how to adequately capture scattering effects for nonlinear transport using other quantum approaches is still a challenge. Rigorous treatment, e.g., the Kohn-Luttinger approach, becomes extremely complicated even for simple models. To simplify, most recent works treat scattering by adding various phenomenological relaxation terms to quantum kinetic equation, which, however, produce unphysical and even self-contradictory results, as discussed in Ref. [40]. In this regard, our work also offers an important guidance for theory development, namely, to achieve consistent physics, any proper treatment of scattering in a theory must at least capture the correct IAD contribution.
Finally, as a by-product, we predict a new phenomenon for ac transport in the high-frequency regime $\tau^{-1}\ll\omega<\Delta$. According to Eq. (19) (and Table 1), in this regime, due to the vanishing double-frequency IAD, the second-harmonic response $j^{(2\omega)}$ acquires a value different from the low-frequency regime and from the rectified component. This leads to a non-Hall and a longitudinal component of the second-harmonic nonlinear current. Using first-principles calculations, we perform an estimation of this phenomenon for antiferromagnetic CuMnAs (see End Matter) and predict a sizable response that can be detected in experiment.